# Robust Beamforming Design and Antenna Selection for Dynamic HRIS-aided Massive MIMO Systems

### v1

@@ III-A Problem Transformation
Since the reflection amplitude matrix ${\bf{\Lambda}}\!=\!{\rm diag}({\bm{\omega}})$ and the mode selection matrix ${\bf{B}}\!=\!{\rm diag}({\bm{\gamma}})$ are tightly coupled, we first define the auxiliary variable ${\bm{\tilde{\omega}}}={\rm diag}({\bf{\Lambda}}{\bf{B}})$ as follows
@@
The binary variables $\{{{\bm{\gamma}},{\bf{a}},{\bf{z}}}\}$ still render the problem intractable and NP-hard. Though the optimal solution can be obtained by the state-of-the-art Gurobi solver embedded with the branch-and-bound (BB) method, the worst-case complexity increases exponentially with the scale of the problem.
To balance the complexity and optimality, we propose a variational reformulation of the binary constraints. Specifically, these constraints can be equivalently transformed into the $l_{2}$ box non-separable constraints with the following lemma.
@@
[…]
$\quad\qquad\quad{\bf{a}}_{i}\neq{\bf{a}}_{j},\forall i\neq j,\{i,j\}\in\mathcal{L},$ (10c)
$\quad\qquad\quad\mu>\mu_{\rm min},$ (10d)
$\quad\qquad\quad{\bm{\tilde{\gamma}}}^{T}{\bf{\tilde{u}}}\!=\!N,{\bf{\tilde{a}}}^{T}{\bf{\tilde{v}}}\!=\!LN_{r},{\bf{\tilde{z}}}^{T}{\bf{\tilde{q}}}\!=\!2^{B}N,$ (10e)
$\quad\qquad\quad||{\bf{\tilde{u}}}||_{2}^{2}\!\leq\!N,||{\bf{\tilde{v}}}||_{2}^{2}\!\leq\!LN_{r},||{\bf{\tilde{q}}}||_{2}^{2}\!\leq\!2^{B}N,$ (10f)
$\quad\qquad\quad{\bf{1}}^{T}{\bf{z}}_{n}=1,\forall n\in\mathcal{N};{\bf{1}}^{T}{\bf{a}}_{i}=1,\forall i\in\mathcal{L},$ (10g)
@@ III-C Amplification Factor
Then we have the following proposition concerning the amplification factor selection:
@@ III-E1 Dynamic mode selection
with ${\bf{E}}_{1}={\mu}^{2}\sigma_{a}^{2}{\rm\widetilde{diag}}({\bf{G}}{\bf{X}}{\bf{G}}^{H})+({\mu}-1)^{2}{\bf{K}}$, ${\bf{E}}_{2}=\mu^{2}({\tilde{p}}{\rm\widetilde{diag}}({\bf{h}}_{r}{\bf{h}}_{r}^{H})+\sigma_{a}^{2}{\bf{I}})$ and ${\bf{e}}=\Re\{2({\mu}-1)({\bf{K}}{\bf{1}}+{\bf{k}})$.
@@ III-E3 Reflection phase selection
Since the optimization problems (P5–7) are quadratic convex problems subject to linear and quadratic constraints, the global optimal solutions can be efficiently obtained via the standard CVX solvers. The penalty $\rho$ is updated with every $T$ iterations to realize a trade-off between the solution accuracy and the computational complexity. Finally, the BCD algorithm converges to a stationary point with a sufficiently large penalty.
@@ IV Simulation
In this section, we conduct numerical simulations to evaluate the performance of the dynamic HRIS-aided communication system. Unless otherwise stated, we set the basic system parameters as follows: $N_{R}=32$, $L=8$, $N=64$, $B=2$, $k_{t}=k_{r}=0.08$. Under a three-dimensional deployment setup, the BS and the HRIS are deployed at $(0,80,5)$m and $(0,50,15)$m, respectively. The users are randomly distributed in a circle with a radius of $8$m and a center of $(0,0,2)$m. The transmit power for the user is $p=10$ dBm. The noise powers are given by $\sigma_{b}^{2}=\sigma_{c}^{2}=\sigma^{2}=-80$dBm. The channel ${\bf{h}}_{d}$ is assumed to be Rayleigh fading, while the HRIS-related channels ${\bf{h}}_{r}$ and ${\bf{G}}$ obey Rician fading with the Rician factor being 0.75. For the large-scale fading, the path loss model $\text{PL}(d)\!=\!\beta_{0}(d)^{-{\alpha}_{p}}$ is considered, where $d$ and ${\alpha}_{p}$ respectively denote the propagation distance and the path loss exponent. $\beta_{0}$ stands for the path loss at the reference distance of 1m and is set as $-30$dB. In particular, we set ${\alpha}_{p}^{RB}=2.2$, ${\alpha}_{p}^{UR}=2.2$ and ${\alpha}_{p}^{UB}=3.5$ for the BS-HRIS channel ${\bf G}$, the HRIS-user channel ${\bf h}_{r}$ and the BS-user channel ${\bf h}_{d}$, respectively. The conventional passive RIS and active RIS schemes are adopted as the benchmarks. Besides, the hybrid RIS with a fixed number of active elements, denoted as “F-HRIS”, is compared to illustrate the superiority of the proposed dynamic HRIS scheme (“D-HRIS”). The non-robust scheme of ignoring the impact of HWIs, represented as “N-HRIS”, is also considered. For a fair comparison, we assume that the total power budgets for different schemes are the same.
@@
The MSE performance versus the transmit power of HRIS $P_{\rm HRIS}$ for different schemes are compared in Fig. 2. In Fig. 2 (a), the proposed D-HRIS achieves the same performance as the active RIS in the high transmit power region. In contrast, the D-HRIS performs better than the active RIS in the lower-power regime. It happens because the transmit power is not sufficient to activate all the elements in the active RIS. The performance gap between the D-HRIS and F-HRIS demonstrates the effectiveness of the dynamic configuration for the active elements. Besides, the MSE performance of passive RIS also increases as the transmit power of HRIS increases, since the total power budget increases correspondingly. In Fig. 2 (b), the ideal D-HRIS limits the MSE performance of the HWI-aware systems to an upper bound. The D-HRIS always achieves better performance than the N-HRIS scheme under various setups of HWIs. It showcases the effectiveness of the robust design under imperfect hardware implementations.
@@ V Conclusion
In this paper, we investigated the channel-aware placement of receive antennas at the BS and active/passive elements at the HRIS under the imperfect hardware implementation. The average MSE minimization problem for the dynamic HRIS-aided massive MIMO system was studied by jointly optimizing the receive antenna selection matrix, the reflection phase coefficients, the reflection amplitude matrix, and the mode selection matrix for the active elements. To tackle the non-convexity and intractability of this problem, we proposed a penalty-based exact BCD algorithm to solve these variables alternately. Numerical simulations showed great superiority of the channel-aware configuration for the receive antennas and active/passive elements in the HRIS-aided massive MIMO system as compared to the conventional benchmark schemes.

### v2

@@ III-A Problem Transformation
Since the reflection amplitude matrix ${\bf{\Lambda}}\!=\!{\rm diag}({\bm{\omega}})$ and the mode selection matrix ${\bf{B}}\!=\!{\rm diag}({\bm{\gamma}})$ are tightly coupled, we first introduce the auxiliary variable ${\bm{\tilde{\omega}}}={\rm diag}({\bf{\Lambda}}{\bf{B}})$, representing the reflection amplification matrix related to those active HRIS elements, as follows
@@
The binary variables $\{{{\bm{\gamma}},{\bf{a}},{\bf{z}}}\}$ still render the problem intractable and NP-hard. Though the optimal solution can be obtained by the state-of-the-art Gurobi solver embedded with the branch-and-bound (BB) method, the worst-case complexity increases exponentially with the scale of the problem. To balance the complexity and optimality, we propose a variational reformulation of the binary constraints. Specifically, these constraints can be equivalently transformed into the $l_{2}$ box non-separable constraints with the following lemma.
@@
s.t. $\mu^{2}{\bm{\gamma}}^{H}({\tilde{p}}{\rm\widetilde{diag}}({\bf{h}}_{r}{\bf{h}}_{r}^{H})\!+\!\sigma_{a}^{2}{\bf{I}}){\bm{\gamma}}\leq P_{\rm HRIS},$ (10b)
${\bf{a}}_{i}\neq{\bf{a}}_{j},\forall i\neq j,\{i,j\}\in\mathcal{L},$ (10c)
$\mu>\mu_{\rm min},$ (10d)
${\bm{\tilde{\gamma}}}^{T}{\bf{\tilde{u}}}\!=\!N,{\bf{\tilde{a}}}^{T}{\bf{\tilde{v}}}\!=\!LN_{r},{\bf{\tilde{z}}}^{T}{\bf{\tilde{q}}}\!=\!2^{B}N,$ (10e)
$||{\bf{\tilde{u}}}||_{2}^{2}\!\leq\!N,||{\bf{\tilde{v}}}||_{2}^{2}\!\leq\!LN_{r},||{\bf{\tilde{q}}}||_{2}^{2}\!\leq\!2^{B}N,$ (10f)
${\bf{1}}^{T}{\bf{z}}_{n}=1,\forall n\in\mathcal{N};{\bf{1}}^{T}{\bf{a}}_{i}=1,\forall i\in\mathcal{L},$ (10g)
@@ III-C Amplification Factor
The $\mu_{\rm ref}$ represents the maximum amplification factor that is constrained by the transmit power allocated to the active elements at the current iteration. Then we have the following proposition concerning the amplification factor selection:
@@ III-E1 Dynamic mode selection
with ${\bf{E}}_{1}={\mu}^{2}\sigma_{a}^{2}{\rm\widetilde{diag}}({\bf{G}}{\bf{X}}{\bf{G}}^{H})+({\mu}-1)^{2}{\bf{K}}$, ${\bf{E}}_{2}=\mu^{2}({\tilde{p}}{\rm\widetilde{diag}}({\bf{h}}_{r}{\bf{h}}_{r}^{H})+\sigma_{a}^{2}{\bf{I}})$ and ${\bf{e}}=\Re\{2({\mu}-1)({\bf{K}}{\bf{1}}+{\bf{k}})\}$.
@@ III-E3 Reflection phase selection
Since the optimization problems (P5–7) are quadratic convex problems subject to linear and quadratic constraints, the global optimal solutions can be efficiently obtained via the standard CVX solvers. The penalty $\rho$ is updated with every $T$ iterations to realize a trade-off between the solution accuracy and the computational complexity.
Finally, the PEBCD algorithm is summarized in Algorithm 1. The proposed algorithm is guaranteed to converge to a stationary solution with a sufficiently large penalty. The computational complexity of Algorithm 1 primarily stems from the update of primal variables in step 3, 4, and 6. In step 3, the computation of matrix inversion and multiplication contribute to a complexity of $O(LN_{R}^{2}+L^{2}N_{R}+L^{3})$. Step 4 involves the matrix multiplication with a complexity of $O(N^{2}N_{R}+NN_{R}^{2})$. The interior point method of the CVX solver has a complex of $O(\log(1/{\epsilon})(2^{3B}N^{3}+L^{3}N_{R}^{3}))$, with $\epsilon$ representing the desired accuracy. Therefore, the total complexity of Algorithm 1 can be estimated as $O(I_{BCD}(\log(1/{\epsilon})(2^{3B}N^{3}+L^{3}N_{R}^{3})+N^{2}N_{R}+NN_{R}^{2}))$, where $I_{BCD}$ denotes the number of iterations for the proposed PEBCD algorithm.
Input : System parameters $N_{R}$, $L$, $N$, $B$, $k_{t}$, $k_{r}$, $\sigma_{b}^{2}$, $\sigma_{c}^{2}$, $p$, and the threshold $\epsilon$.
Output : ${\bf{w}}^{\star}$,${\bf{A}}^{\star},{\bf{B}}^{\star}$, ${\bm{\Phi}}^{\star},{\bm{\Lambda}}^{\star}$.
1
2Initialize $({\bf{w}}^{0},{{\mu}}^{0},{\bf{v}}^{0},{\bf{u}}^{0},{\bf{q}}^{0},{\bm{\gamma}}^{0},{\bf{a}}^{0},{\bf{z}}^{0})$, and set the initial penalty ${\rho}$.
3 repeat
4
5      Update the receive beamforming vector ${\bf{w}}^{t}$ by (13),
6
7      Update the amplification factor $\mu^{t}$ by (23),
8
9      Update the auxiliary variables $\{{\bf{u}}^{t},{\bf{v}}^{t},{\bf{q}}^{t}\}$ by (24),
10
11      Update the selection variables $\{{\bm{\gamma}}^{t},{\bf{a}}^{t},{\bf{z}}^{t}\}$ via the CVX solver by solving (III-E1), (III-E2) and (III-E3),
12
13      If $(t\mod T)=0$
14         Increase the penalty $\rho^{t}$ ,
15       End
16until  Convergence ;
Recover ${\bf{A}}^{\star}$, ${\bf{B}}^{\star}$, ${\bm{\Phi}}^{\star}$, ${\bm{\Lambda}}^{\star}$ from ${{\mu}}^{\star}$, ${\bf{v}}^{\star}$, ${\bf{u}}^{\star}$, ${\bf{q}}^{\star}$, ${\bm{\gamma}}^{\star}$, ${\bf{a}}^{\star}$, ${\bf{z}}^{\star}$.
Algorithm 1 Proposed PEBCD Algorithm
@@ IV Simulation
In this section, we conduct numerical simulations to evaluate the performance of the dynamic HRIS-aided communication system. Unless otherwise stated, we set the basic system parameters as follows: $N_{R}=32$, $L=8$, $N=64$, $B=2$, $k_{t}=k_{r}=0.08$ [16]. Under a three-dimensional deployment setup, the BS and the HRIS are deployed at $(0,80,5)$m and $(50,50,15)$m, respectively. The user is located at $(0,0,2)$m. The transmit power for the user is $p=10$ dBm. The noise powers are given by $\sigma_{b}^{2}=\sigma_{c}^{2}=\sigma^{2}=-80$dBm. The channel ${\bf{h}}_{d}$ is assumed to be Rayleigh fading, while the HRIS-related channels ${\bf{h}}_{r}$ and ${\bf{G}}$ obey Rician fading with the Rician factor being 0.75. For the large-scale fading, the path loss model $\text{PL}(d)\!=\!\beta_{0}(d)^{-{\alpha}_{p}}$ is considered, where $d$ and ${\alpha}_{p}$ respectively denote the propagation distance and the path loss exponent. $\beta_{0}$ stands for the path loss at the reference distance of 1m and is set as $-30$dB. In particular, we set ${\alpha}_{p}^{RB}=2.2$, ${\alpha}_{p}^{UR}=2.2$ and ${\alpha}_{p}^{UB}=3.5$ for the BS-HRIS channel ${\bf G}$, the HRIS-user channel ${\bf h}_{r}$ and the BS-user channel ${\bf h}_{d}$, respectively. The conventional passive RIS and active RIS schemes are adopted as the benchmarks. Besides, the hybrid RIS with a fixed number of active elements, denoted as “F-HRIS”, is compared to illustrate the superiority of the proposed dynamic HRIS scheme (“D-HRIS”). The non-robust scheme of ignoring the impact of HWIs, represented as “N-HRIS”, is also considered. “D-HRIS w/o AS” denotes the dynamic HRIS-aided MISO communication system with fixed BS antenna selection. For a fair comparison, we assume that the total power budgets for different schemes are the same.
@@
Figure 3: Robust versus nonrobust design under the transmit power of HRIS.
The MSE performance versus the transmit power of HRIS $P_{\rm HRIS}$ for different schemes are compared in Fig. 2. In Fig. 2, the performance of the passive RIS scheme improves with an increase in the transmit power of the HRIS, denoted as $P_{HRIS}$. This improvement is due to the corresponding increase in the transmit power at the BS for passive RIS to ensure a fair comparison. The proposed D-HRIS achieves the same performance as the active RIS in the high transmit power region. In contrast, the D-HRIS performs better than the active RIS in the lower-power regime. It happens because the transmit power is not sufficient to activate all the elements in the active RIS. The performance gap between the D-HRIS and F-HRIS demonstrates the effectiveness of the dynamic configuration for the active elements. Besides, the MSE performance of passive RIS also increases as the transmit power of HRIS increases, since the total power budget increases correspondingly. The performance gap between the “D-HRIS” and “D-HRIS w/o AS” schemes demonstrates the necessity of BS antenna selection, as it can effectively combat multipath fading by adjusting the antennas. In Fig. 3, the ideal D-HRIS limits the MSE performance of the HWI-aware systems to an upper bound. The D-HRIS always achieves better performance than the N-HRIS scheme under various setups of HWIs. It showcases the effectiveness of the robust design under imperfect hardware implementations.
@@ V Conclusion
In this paper, we investigated the channel-aware placement of receive antennas at the BS and active/passive elements at the HRIS under the imperfect hardware implementation. The average MSE minimization problem for the dynamic HRIS-aided MISO system was studied by jointly optimizing the receive antenna selection matrix, the reflection phase coefficients, the reflection amplitude matrix, and the mode selection matrix for the active elements. To tackle the non-convexity and intractability of this problem, we proposed a PEBCD algorithm to solve these variables alternately. Numerical simulations showed great superiority of the channel-aware configuration for the receive antennas and active/passive elements in the HRIS-aided MISO system as compared to the conventional benchmark schemes.